OpenStudy (anonymous):

Multiply. 8 x 2 3/4 =

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

probably best to turn \[2\tfrac{3}{4}\] into an improper fraction

OpenStudy (anonymous):

Improper fraction: 12/4

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

\[\frac{12}{4}=3\]

OpenStudy (anonymous):

Oh. Okay thank you so much!

OpenStudy (anonymous):

hold on we are not done at all still have to convert \[2\tfrac{3}{4}\] into and improper fraction

OpenStudy (anonymous):

do you know how to do it?

OpenStudy (anonymous):

12/4 = 3

OpenStudy (anonymous):

yes but you do not have 3, you have \[2\tfrac{3}{4}\] so that is wrong

OpenStudy (anonymous):

\[2\tfrac{3}{4}=\frac{2\times 4+3}{4}\]